Jenna has just enough money to rent a canoe for 1.5 hours. How far out on a lake can she paddle and return on time if she paddles out at 2 kph and back at 4 kph?

Answer by stanbon(75887)

Jenna has just enough money to rent a canoe for 1.5 hours. How far out on a lake can she paddle and return on time if she paddles out at 2 kph and back at 4 kph?
----
Out DATA:
rate = 2 km/h ; time = x hrs ; distance = r*t = 2x km
-----
Back DATA:
rate = 4 km/h ; time = 1.5-x hrs ; distance = 4(1.5-x) km
-----
Equation:
distance out = distance back
2x = 4(1.5-x)
2x = 6-4x
6x = 6
x = 1 hr
-----
Ans: distance out = 2x = 2*1 = 2 km
--------------
Cheers,
Stan H.
----------------